OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

Remember PEMDAS 4(6 + 7) = 4 • 6 + x • 7 Parenthesis 4(13)=4 • 6 + x • 7 Multiply 52=24 + 7x Now we need to isolate x Subtract 24 on both sides 52=24 + 7x -24 -24 28=7x Divide by 7 on both sides 28=7x /7 /7 4=x or x=4